ganeshie8 (ganeshie8):

Find \(2000\) th term of below sequence \[2,3,5,6,7,8,10,\ldots\] (the sequence contains positive integers other than perfect squares)

Nnesha (nnesha):

\(\huge\color{white}{{ *}}\)

OpenStudy (nincompoop):

993

OpenStudy (nincompoop):

my bad

OpenStudy (nincompoop):

1993

OpenStudy (nincompoop):

sequence of a number that are not squares

ganeshie8 (ganeshie8):

thats very close ! but it has to be greater than 2000 for sure... as the 2000th term of positive integer sequence(1,2,3,4,...) itself is 2000. i am looking for answer and a neat way to justify the answer :)

OpenStudy (nincompoop):

the 2000th term is 2000? I do not think so. We started with 2 and skipped 4 and 9 that is down to 10 - 3 sequence

ganeshie8 (ganeshie8):

2000th term of below sequence is 2000 : \[1,2,3,4,5, \ldots\] so the 2000th term of the sequence in main question has to be greater than 2000

OpenStudy (nincompoop):

ohh hahahha

OpenStudy (nincompoop):

right

OpenStudy (nincompoop):

can I just give you the pseudo-code in Python or C++? laughing out loud

ganeshie8 (ganeshie8):

lol it can be considered a sledgehammer in the context of this problem :P

OpenStudy (anonymous):

2044?

OpenStudy (anonymous):

till 2000.. 44 squares are missing from the series last of which is 1936.. so 2044

ganeshie8 (ganeshie8):

thats very close too.. but no!

OpenStudy (anonymous):

then 2045?

ganeshie8 (ganeshie8):

2045 is right ! :)

OpenStudy (nincompoop):

how do I interpret this? \(a_n = \lfloor \sqrt{n}+\frac{1}{2})\rfloor +n\)

ganeshie8 (ganeshie8):

that looks interesting xD is that the general term of sequence ?

OpenStudy (nincompoop):

it is, apparently

OpenStudy (nincompoop):

don't mind me, I am crashing from the caffeine-induced waking state

ganeshie8 (ganeshie8):

that actually works! http://www.wolframalpha.com/input/?i=floor%28sqrt%282000%29%2B1%2F2%29%2B2000

OpenStudy (anonymous):

2045? 1935:1979 1956:2000 1980:2024 1981:2026 2000:2045

OpenStudy (ikram002p):

i thought of it like this sequence of perfect square up to some value http://www.wolframalpha.com/input/?i=n%5E2+from+n%3D1+to+n%3D50 so we can note that 2000-44=1956 so the 1956 term is 2000 now we keep moving like this 2001.. to 2024 ( the 1980 term ) now we need other 20 terms :P from 2026 .... to 2045 so i would say 2045

mathslover (mathslover):

Suscribing to the post :P

mathslover (mathslover):

Oops *Subscribing

jaynator495 (jaynator495):

oh hia der...

OpenStudy (kainui):

Yeah I like @ikram002p 's way, that's how I was gonna do it haha.